Suppose you are thinking of purchasing the stock of TTYL Widget, Inc.  You expect it to pay a $3.00 dividend in one year.  You believe you can sell the stock for $21.00 at that time.  You require a return of 15% on investments of this risk.  What is the maximum you would be willing to pay?

Suppose you are thinking of purchasing the stock of TTYL Widget, Inc.  In addition to the dividend and price from year one you expect it to pay a $3.30 dividend in two years.  You believe you can sell the stock for $23.10 at that time.  You require a return of 15% on investments of this risk.  What is the maximum you would be willing to pay?

Suppose you are thinking of purchasing the stock of ABC Widget, Inc.  In addition to the dividend and price from year one and two you expect it to pay a $3.63 dividend in three years.  You believe you can sell the stock for $25.41 at that time.  You require a return of 15% on investments of this risk.  What is the maximum you would be willing to pay?

Suppose a stock is expected to pay a $1.25 dividend every quarter and the required return is 8% with quarterly compounding. What is the price?

Suppose Pizza just paid a dividend of $.50. It is expected to increase its dividend by 2% per year. If the market requires a return of 15% on assets of this risk, how much should the stock be selling for?
Answer #1

1.
=(3+21)/(1+0.15)=20.8695652173913

2.
=3/1.15+3.3/1.15^2+23.1/1.15^2=22.5708884688091

3.
=3/1.15+3.3/1.15^2+3.63/1.15^3+25.41/1.15^3=24.1982411440783

4.
=1.25/(8%/4)=62.5

5.
=0.5*(1+2%)/(15%-2%)=3.92307692307692